The ratios of stars to squares is 2 to 3. if there were 6 stars, how many squares would there be what is the correct answer' in sentences
1
Expert's answer
2020-09-30T12:07:33-0400

Ratio of stars and square = 2/3 , it means

=> {no of stars/no of squares} = 2/3 .................(1)

Now as given in question, No of stars =6.

=> 6/ (No of squares) = 2/3

=> No of squares = 6x3/2 = 9

Hence "No of Squares are 9"